# SmartBench: Is Your LLM Truly a Good Chinese Smartphone Assistant?

### v1

@@ Abstract
Large Language Models (LLMs) have become integral to daily life, especially advancing as intelligent assistants through on-device deployment on smartphones. However, existing LLM evaluation benchmarks predominantly focus on objective tasks like mathematics and coding in English, which do not necessarily reflect the practical use cases of on-device LLMs in real-world mobile scenarios, especially for Chinese users. To address these gaps, we introduce SmartBench, the first benchmark designed to evaluate the capabilities of on-device LLMs in Chinese mobile contexts. We analyze functionalities provided by representative smartphone manufacturers and divide them into five categories: text summarization, text Q&A, information extraction, content creation, and notification management, further detailed into 20 specific tasks. For each task, we construct high-quality datasets comprising 50 to 200 question-answer pairs that reflect everyday mobile interactions, and we develop automated evaluation criteria tailored for these tasks. We conduct comprehensive evaluations of on-device LLMs and MLLMs using SmartBench and also assess their performance after quantized deployment on real smartphone NPUs. Our contributions provide a standardized framework for evaluating on-device LLMs in Chinese, promoting further development and optimization in this critical area. Code and data will be available at https://github.com/Lucky-Lance/SmartBench.
@@
Xudong Lu*1,2, Haohao Gao*1, Renshou Wu*$\dagger$1, Shuai Ren1, Xiaoxin Chen1, Hongsheng ${}^{\textrm{\Letter}}$, Fangyuan ${}^{\textrm{\Letter}}$ 1vivo AI Lab 2CUHK MMLab Equal contribution ${}^{\textrm{\Letter}}$orresponding author Project lead {luxudong@link,hsli@ee}.cuhk.edu.hk{wurenshou,lifangyuan}@vivo.com
@@ 2.4 LLM Agent on Smartphones
Currently, there is another type of task on mobile phones that helps solve real-world tasks, called mobile agents Wang et al. (2024a); Zhang et al. (2023a); Chai et al. (2024). These tasks often involve executing multi-step commands on the phone based on user instructions Zhang et al. (2024). In contrast, Smartbench focuses on the functionality of on-device LLMs for handling common daily tasks in a single step, without the need for planning action trajectories or calling external APIs.
@@ 3.1 Data Composition
Category Task BlueLM-3B InternVL2.5-4B MiniCPM3-4B Qwen2.5-3B Qwen2-VL-2B
Text Summarization Document Summ 7.56 6.89 7.40 7.21 4.37
Call Summ 7.22 5.43 6.88 6.35 3.48
Recording Summ 7.63 6.38 7.45 7.07 4.17
Meeting Summ 7.09 6.23 6.98 6.67 3.75
Text Q&A Document Q&A 9.37 9.36 8.39 9.34 9.15
Retrieval Q&A 5.89 5.81 6.76 6.25 4.77
Personal Q&A 9.36 8.89 8.87 9.39 8.83
Content Creation Text Polishing 7.55 6.17 7.53 7.42 6.19
Text Continuation 7.45 6.89 7.52 7.72 5.96
Text Abbreviation 8.23 7.43 8.17 8.51 7.51
Text Expansion 7.44 8.04 8.74 8.07 6.04
Text Creation 6.93 6.16 6.89 6.68 5.26
Text Formatting 6.03 5.10 6.80 3.69 1.20
Instant Reply 6.70 5.90 6.28 6.44 3.14
Text Correction 3.69 2.46 3.24 2.38 1.17
Information Extraction Entity Extraction 7.82 8.13 7.58 6.35 5.00
Relation Extraction 5.55 3.58 4.15 3.54 3.04
Event Extraction 6.79 7.09 6.20 6.75 4.66
Notification Management Message Summ 7.45 7.29 8.08 7.86 5.90
Notification Sorting 4.78 4.19 4.85 4.51 2.14
AVG 7.03 6.37 6.94 6.61 4.79
Table 3: Evaluation results using GPT-4 Turbo (gpt-4-turbo-04-09) as the judge LLM. We compare BlueLM-3B, InternVL2.5-4B, MiniCPM3-4B, Qwen2.5-3B, and Qwen2-VL-2B on the whole SmartBench with BF16 precision. BlueLM-3B achieves the best overall performance, with the highest average score and the most bolded metrics (10 out of 20 tasks). The scores assessed by Qwen-Max as the judge LLM are also provided in Tab. 7 in the Appendix.
@@ 4.1 BF16 Precision Evaluation
In this subsection, we evaluate representative on-device LLMs/MLLMs on SmartBench (BF16 parameter precision). We select BlueLM-3B Lu et al. (2024b), InternVL2.5-4B Chen et al. (2024), MiniCPM3-4B Hu et al. (2024a), Qwen2.5-3B Yang et al. (2024b), and Qwen2-VL-2B Wang et al. (2024b). GPT-4 Turbo (gpt-4-turbo-04-09) is utilized as the judge LLM. The results are summarized in Tab. 3, where BlueLM-3B achieves the highest average score. Additionally, we can observe the following trends from the table:
@@
For a more comprehensive evaluation, we also present the scores assessed by Qwen-Max (qwen-max-longcontext) as the judging LLM in Tab. 7 in the Appendix. It can be observed that although there are slight differences in the average scores, both Qwen-Max and GPT-4 Turbo rank the models in the same order. This demonstrates the robustness of our LLM-as-a-Judge approach.
Remark: We evaluate MLLMs on SmartBench because, in on-device deployment scenarios on real smartphones, memory limitations often prevent us from deploying both an LLM and an MLLM on the device. Consequently, this on-device model must simultaneously handle both pure language tasks and multimodal tasks effectively.
@@ 4.2 INT4 Precision Evaluation on NPU
On-device LLMs are often deployed on the smartphone’s Neural Processing Unit (NPU) to leverage its specialized parallel computational capabilities. In our experiment, we deploy the BlueLM-3B and Qwen2.5-3B models on the NPU of the vivo iQOO 12 smartphone equipped with the Snapdragon 8 Gen 3 SoC. To be specific, we quantize the models to W4A16 using the Qualcomm QNN SDK [^1] https://www.qualcomm.com/developer/software/neural-processing-sdk-for-ai. Due to the inference speed limitations on the mobile NPU, we select 50 questions per task for inference.
The results are shown in Tab. 4. We present the scores for each task (BF16 and INT4) and the capability retention of the INT4 models. For most tasks, the quantized models retain over 80% of their original capabilities, with an overall average retention rate of approximately 90%. Additionally, we provide the evaluation results using Qwen-Max as the judge LLM in Tab. 8 in the Appendix.
@@ 4.3 Human Test
In SmartBench, we meticulously design evaluation dimensions and scoring standards for each task/category. To establish a baseline, we compare our evaluation prompts with those used in MT-Bench. We randomly select 20 questions for each task, with each question containing responses from 5 on-device models. This results in a total of 20$\times$20$\times$5=2000 samples. We conduct human ranking and calculate the Pearson correlation with the judge LLM ranking (our prompt versus MT-Bench prompt), and the results are shown in Tab. 5. Our designed prompt excels in all categories.
@@ 5 Conclusion
In this paper, we present SmartBench, the first benchmark specifically designed to evaluate the capabilities of on-device LLMs in Chinese mobile contexts. By analyzing functionalities offered by leading smartphone manufacturers, we create a standardized framework divided into five key categories and 20 specific tasks, complete with high-quality datasets and tailored evaluation criteria. Our comprehensive evaluations of on-device LLMs and MLLMs using SmartBench highlight the strengths and weaknesses of current models in real-world mobile scenarios. This work fills a critical gap in benchmarking tools for Chinese users, promoting further development and optimization of on-device LLMs in practical mobile applications. We hope our work can contribute to the progress of this area of research.
@@
In this paper, we provide SmartBench, the first benchmark designed to evaluate the capabilities of on-device LLMs in Chinese mobile contexts. Our work still has some limitations: 1) With the advancement of technology, the functions of on-device LLMs will continually evolve. Our investigation only covers up to December 2024. We will continue to update the dataset in line with the release of new features. 2) We have developed SmartBench specifically for the usage scenarios of Chinese users. The usage habits and methods of smartphone users may vary significantly across different countries. Moving forward, we will continue to support multiple languages.
@@ A.2 More Evaluation Results
Text Summarization Document Summ 7.20 6.98 6.95 7.09 4.74
@@
Content Creation Text Polishing 6.93 5.78 6.90 6.86 5.91
@@
Information Extraction Entity Extraction 7.77 7.40 7.44 6.21 5.00
@@
Notification Management Message Summ 6.92 6.96 7.47 7.62 5.64
@@ A.3 Details of Human Annotators
A.4 More Example Cases
@@
A.5 English Translation of Pictures in the Paper

### v2

@@ Abstract
Large Language Models (LLMs) have become integral to daily life, especially advancing as intelligent assistants through on-device deployment on smartphones. However, existing LLM evaluation benchmarks predominantly focus on objective tasks like mathematics and coding in English, which do not necessarily reflect the practical use cases of on-device LLMs in real-world mobile scenarios, especially for Chinese users. To address these gaps, we introduce SmartBench, the first benchmark designed to evaluate the capabilities of on-device LLMs in Chinese mobile contexts. We analyze functionalities provided by representative smartphone manufacturers and divide them into five categories: text summarization, text Q&A, information extraction, content creation, and notification management, further detailed into 20 specific tasks. For each task, we construct high-quality datasets comprising 50 to 200 question-answer pairs that reflect everyday mobile interactions, and we develop automated evaluation criteria tailored for these tasks. We conduct comprehensive evaluations of on-device LLMs and MLLMs using SmartBench and also assess their performance after quantized deployment on real smartphone NPUs. Our contributions provide a standardized framework for evaluating on-device LLMs in Chinese, promoting further development and optimization in this critical area. Code and data will be available at https://github.com/vivo-ai-lab/SmartBench.
@@
[…]
@@ 2.4 LLM Agent on Smartphones
There is another type of task on mobile phones that helps solve real-world tasks, called mobile agents Wang et al. (2024a); Zhang et al. (2023a); Chai et al. (2024); Rawles et al. (2024). These tasks often involve executing multi-step commands on the phone based on user instructions Zhang et al. (2024). In contrast, Smartbench focuses on the functionality of on-device LLMs for handling common daily tasks in a single step, without planning action trajectories or calling external APIs.
@@ 3.1 Data Composition
Category Task GPT-4o BlueLM-3B InternVL2.5-4B MiniCPM3-4B Qwen2.5-3B Qwen2-VL-2B
Text
Summarization
Document Summ 7.05 7.56 6.89 7.40 7.21 4.37
Call Summ 7.03 7.22 5.43 6.88 6.35 3.48
Recording Summ 7.78 7.63 6.38 7.45 7.07 4.17
Meeting Summ 7.73 7.09 6.23 6.98 6.67 3.75
Text Q&A Document Q&A 8.83 9.37 9.36 8.39 9.34 9.15
Retrieval Q&A 6.91 5.89 5.81 6.76 6.25 4.77
Personal Q&A 9.78 9.36 8.89 8.87 9.39 8.83
Content
Creation
Text Polishing 7.49 7.55 6.17 7.53 7.42 6.19
Text Continuation 7.63 7.45 6.89 7.52 7.72 5.96
Text Abbreviation 7.49 8.23 7.43 8.17 8.51 7.51
Text Expansion 8.42 7.44 8.04 8.74 8.07 6.04
Text Creation 7.55 6.93 6.16 6.89 6.68 5.26
Text Formatting 6.15 6.03 5.10 6.80 3.69 1.20
Instant Reply 7.62 6.70 5.90 6.28 6.44 3.14
Text Correction 7.03 3.69 2.46 3.24 2.38 1.17
Information
Extraction
Entity Extraction 8.36 7.82 8.13 7.58 6.35 5.00
Relation Extraction 3.54 5.55 3.58 4.15 3.54 3.04
Event Extraction 7.86 6.79 7.09 6.20 6.75 4.66
Notification
Management
Message Summ 7.24 7.45 7.29 8.08 7.86 5.90
Notification Sorting 5.97 4.78 4.19 4.85 4.51 2.14
AVG 7.37 7.03 6.37 6.94 6.61 4.79
Table 3: Evaluation results using GPT-4 Turbo (gpt-4-turbo-04-09) as the judge LLM. We compare BlueLM-3B, InternVL2.5-4B, MiniCPM3-4B, Qwen2.5-3B, and Qwen2-VL-2B (on-device models) on the whole SmartBench in BF16 precision with GPT-4o (on-cloud model) for reference. The scores assessed by Qwen-Max as the judge LLM are also provided in Tab. 8 in the Appendix.
@@ 3.3 Data Screening
To be specific, we implement a five-point scoring system (1–5) for each criterion, where 1 = Unacceptable, 2 = Poor, 3 = Fair, 4 = Good, and 5 = Very Good. Human experts score each item across all criteria, and we retain only those with an average score of $\geq$ 3.5, reducing the original 30k dataset to roughly 3k high-quality entries. For items scoring between 3.5 and 4, we perform manual re-labeling to guarantee accuracy and alignment with human judgment. All items are further refined and subjected to dual-layer verification, ensuring rigorous quality control throughout the dataset.
@@ 4.1 BF16 Precision Evaluation
In this subsection, we evaluate representative on-device LLMs/MLLMs on SmartBench (BF16 parameter precision). We select BlueLM-3B Lu et al. (2024b), InternVL2.5-4B Chen et al. (2024), MiniCPM3-4B Hu et al. (2024a), Qwen2.5-3B Yang et al. (2024b), and Qwen2-VL-2B Wang et al. (2024b). GPT-4 Turbo (gpt-4-turbo-04-09) is utilized as the judge LLM. We also provide the scores of GPT-4o to compare on-cloud models with on-device models. The results are summarized in Tab. 3, where BlueLM-3B achieves the highest average score among on-device models. Additionally, we can observe the following trends from the table:
@@
#Params Context Length Prefilling Speed (token/s) Output Speed (token/s) Power (W)
BlueLM-3B 2.7B 2048 930.9 27.1 6.4
Qwen2.5-3B 3.1B 2048 873.4 24.9 6.8
Table 5: Inference speed and power usage of BlueLM-3B and Qwen2.5-3B on iQOO 12 with Qualcomm QNN SDK. Due to its larger parameter size, Qwen2.5-3B exhibits slower inference speed and higher power consumption.
@@
3) On-device models still exhibit notable performance gaps compared to the on-cloud model, particularly on the Text Correction task, where they achieve only about half the score of GPT-4o. This suggests that enhancing the reasoning capability of on-device models remains an important area.
4) For a more comprehensive evaluation, we also present the scores assessed by Qwen-Max (qwen-max-longcontext) as the judging LLM in Tab. 8 in the Appendix. It can be observed that although there are slight differences in the average scores, both Qwen-Max and GPT-4 Turbo rank the models in the same order. This demonstrates the robustness of our LLM-as-a-Judge approach.
Remark: We evaluate MLLMs on SmartBench because in on-device deployment scenarios on real smartphones, memory limitations often prevent us from deploying both an LLM and an MLLM on the device. Consequently, this on-device model must simultaneously handle both pure language tasks and multimodal tasks effectively.
@@ 4.2 INT4 Precision Evaluation on NPU
On-device LLMs are often deployed on the smartphone’s Neural Processing Unit (NPU) to leverage its specialized parallel computational capabilities. In our experiment, we deploy the BlueLM-3B and Qwen2.5-3B models on the NPU of the vivo iQOO 12 smartphone equipped with the Snapdragon 8 Gen 3 SoC. To be specific, we quantize the models to W4A16 using the Qualcomm QNN SDK. Due to the inference speed limitations on the mobile NPU, we select 50 questions per task for inference. The results are shown in Tab. 4. We present the scores for each task (BF16 and INT4) and the capability retention of the INT4 models. Additionally, we provide the evaluation results using Qwen-Max as the judge LLM in Tab. 9 in the Appendix.
1) For most tasks, the quantized models retain over 80% of their original capabilities, with an overall average retention rate of approximately 90%.
2) Although the models can achieve, on average, 90% of the original score on the edge side, they may still generate incorrect responses after quantization. We provide two failure cases of BlueLM-3B in Sec. A.5, where the model exhibits fluency degradation and reduced understanding capability.
3) To offer deeper insights into real-world edge-side deployment, we report the prefilling speed, output token generation speed, and power consumption on the iQOO 12 smartphone using the Qualcomm QNN SDK, as shown in Tab. 5.
@@ 4.3 Human Test
[…]
@@ 5 Conclusion
In this paper, we present SmartBench, the first benchmark designed to evaluate the capabilities of on-device LLMs in Chinese mobile contexts. By analyzing functionalities offered by leading smartphone manufacturers, we create a standardized framework divided into five key categories and 20 specific tasks, complete with high-quality datasets and tailored evaluation criteria. Our comprehensive evaluations of on-device LLMs and MLLMs using SmartBench highlight the strengths and weaknesses of current models in real-world mobile scenarios. This work fills a critical gap in benchmarking tools for Chinese users, promoting further development and optimization of on-device LLMs in practical mobile applications.
@@
In this paper, we provide SmartBench, the first benchmark designed to evaluate the capabilities of on-device LLMs in Chinese mobile contexts. Our work still has some limitations: 1) With the advancement of technology, the functions of on-device LLMs will continually evolve. Our investigation only covers up to December 2024. We will continue to update the dataset in line with the release of new features. 2) We have developed SmartBench specifically for the usage scenarios of Chinese users. The usage habits and methods of smartphone users may vary significantly across different countries. Moving forward, we will continue to support multiple languages. 3) The current benchmark focuses on the text modality, whereas mobile applications may also involve vision and audio modalities (e.g., camera input and voice recognition). We plan to incorporate these additional modalities in future versions.
@@ A.2 More Evaluation Results
Text
Summarization
Document Summ 7.20 6.98 6.95 7.09 4.74
@@
Content
Creation
Text Polishing 6.93 5.78 6.90 6.86 5.91
@@
Information
Extraction
Entity Extraction 7.77 7.40 7.44 6.21 5.00
@@
Notification
Management
Message Summ 6.92 6.96 7.47 7.62 5.64
@@ A.3 Details of Human Annotators
A.4 Comparison with Traditional Benchmarks
Traditional benchmarks (e.g., MMLU) mainly evaluate the model’s objective knowledge, while SmartBench focuses on subjective data aligned with end-side smartphone scenarios, assessing the degree of alignment with human preferences. In practice, due to memory and storage constraints on mobile devices, only a single MLLM can be deployed on the device. We consider the training process from LLM to MLLM. For example, InternVL2.5-4B is trained from Qwen2.5-3B-Instruct. During this process, objective knowledge tends to be preserved, but the subjective (human alignment) performance often degrades.
MMLU SmartBench (GPT-4) SmartBench (Qwen-Max)
Qwen2.5-3B 66.31 6.61 6.49
InternVL2.5-4B 68.35 6.37 6.23
MMLU measures objective knowledge, while SmartBench evaluates subjective alignment. We test SmartBench performance using both GPT-4 and Qwen-Max as evaluators. As can be seen, during the training process from LLM to multimodal models, MMLU performance improves due to the injection of new knowledge. However, the subjective performance represented by SmartBench decreases.
A.5 Failure Cases on NPU
We here provide two failure cases of BlueLM-3B after quantized deployment on the NPU.
Figure 5: Failure case after quantization on the NPU.
Figure 6: Failure case after quantization on the NPU.
A.6 Example Prompt for Generating Task Queries
Figure 7: Prompt for generating mobile notification sorting pairs (en).
A.7 More Example Cases
@@
A.8 English Translation of Pictures in the Paper